OpenStudy (anonymous):

What is the probability that a die is rolled six times and gives the numbers 1-6 in order?

OpenStudy (mathstudent55):

Each roll of an unbiased die is an independent event. The total number of outcomes of six consecutive rolls is 6^6. The probability of one specific outcome is 1/6^6

OpenStudy (zehanz):

The outcomes are not connected. So the probability is 1/6, every time. You need to multiply the probabilities.

OpenStudy (anonymous):

@mathstudent55 is it written like this? 

OpenStudy (zehanz):

No, it's \[\left( \frac{ 1 }{ 6 } \right)^6=\frac{ 1 }{ 6^6 }=\frac{ 1 }{ 46656 }\]

OpenStudy (anonymous):

I see now. Thank you!

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (zehanz):

yw!